OpenStudy (anonymous):

If f'(x) = 3 cos (x/4) and f(pi) = 9 sqrt(2), find f(x) using integration.

OpenStudy (anonymous):

can you write just the problem out no words?

OpenStudy (anonymous):

http://www.wolframalpha.com use this to help you

OpenStudy (anonymous):

@DragonSage did it help?

OpenStudy (anonymous):

nope it gave part of the answer this is the question: f(x) = ∫4 sin x/4 , f(π) = 9√2 The answer is 3√2 + 12 sin (x/4) And I want to know to how to figure it out.

OpenStudy (anonymous):

not sure on this one sorry

OpenStudy (anonymous):

@linda3

OpenStudy (anonymous):

@Data_LG2

OpenStudy (anonymous):

@BTaylor

OpenStudy (anonymous):

anyone help? i cant figure this out

OpenStudy (anonymous):

@ganeshie8

OpenStudy (btaylor):

You have \(f'(x)=\cos \left( \frac{x}{4} \right)\) and have the point \( (\pi,9\sqrt{2})\). \(f(x)\) is going to be the integral of \(f'(x)\). Can you do this part? \[f(x) = \int\limits \cos \left(\frac{x}{4} \right) dx\] Don't forget +C!

OpenStudy (anonymous):

\[f(x) = -4\sin(x/8)\]

OpenStudy (btaylor):

Oops, I forgot the 3 cos... part. Sorry. @DragonSage Can you show your work so we can help you integrate this correctly?

OpenStudy (anonymous):

\[f(x) = 3 \cos \left(\frac{x^2}{8} \right) + c\]

OpenStudy (btaylor):

not quite.

OpenStudy (btaylor):

You will end up with 3(something) sin(something) + C

OpenStudy (anonymous):

\[-3\sin((x^2/8))\]

OpenStudy (btaylor):

almost. It will be a positive sin, since when you take the derivative of + sin you get + cos. Have you learned about substitution?

OpenStudy (anonymous):

isn't the derivative of cos x = -sinx ?

OpenStudy (btaylor):

Yes, but the ANTIderivative of cos x = sin x.

OpenStudy (anonymous):

right sorry I forgot

OpenStudy (anonymous):

so it is \[3 \sin (x^2/8)\] , right?

OpenStudy (btaylor):

No. You don't square the x. cos(u) is the main function. You have to multiply by 4 to compensate.

OpenStudy (anonymous):

oh so that makes it \[12 \sin (x/8)\]

OpenStudy (anonymous):

where does \[3\sqrt2\] come in?